OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

just substitute

OpenStudy (anonymous):

\[x - y = 16\] \[2x + y = 2\] 3x = 18 x=6

OpenStudy (anonymous):

its like: x-y=16 x=16+y put in 2x+y=2 for x 2(16+y)+y=2 32+2y+y=2 3y=2-32 y=-30/3 y=-10 plug this in x-y=16 & solve 4 x to get the value; x-(-10)=16 x+10=16 x=6 so, (6,-10)

OpenStudy (anonymous):

YOLO!!

OpenStudy (accidentalaichan):

^YOLO isn't going to help you on a *ahem* written response *ahem* test.